# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v4

@@ 5 Conclusion
The discussion surrounding zonal and nodal pricing has a long history. The European electricity market has employed zonal pricing for many years, often with large, nationwide price zones. However, the ongoing energy transition requires a reevaluation of price zones. Based on recent data released by ENTSO-E in the context of the EU Bidding Zone Review process, we compare various zonal and nodal pricing rules for the German power market. Our findings indicate that, in terms of system costs and price signals, a nodal allocation rule with non-uniform pricing leads to the lowest total costs. Costly redispatch is mitigated, and the average prices increase only slightly. Additionally, pricing rules such as Join pricing require very low side-payments, while efficiently signaling transmission bottlenecks. In contrast, the proposed zonal configurations only differ marginally regarding system costs and the differences in average prices across zones are low. Also, the effect of zonal splits on price standard deviations is low, and we do not find evidence for significant reductions in redispatch costs resulting from such splits. The study does not say that there cannot be congestion on cross-zonal links leading to significant price differences between zones at certain times. However, in the BZR dataset that was intended to inform ACER based on realistic demand and supply scenarios (the 2009 climate year), we do not find significant price differences. Fears of substantial price differences after a zonal split might be overrated based on this comprehensive and up-to-date dataset. Importantly, we could provide a fair comparison with nodal pricing rules and show that the welfare gains are around 9-10%, depending on the redispatch model.
Acknowledgements
We would like to thank Karsten Neuhoff, Maximilian Rinck, and Anthony Papavasiliou for helfpul comments on earlier versions of the manuscript. Mistakes are of course all ours. This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. BI 1057/10.
Appendix A Matching of Generators to Grid Locations
@@
As discussed in Section 3, we first identify each thermal unit’s broad plant type (e.g., hard coal, gas, etc.) by matching IDs with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur. Next, we need to distribute these plants to the sub-categories provided in the seller data. For example, all hard coal units must be distributed between Hard Coal old1, Hard Coal old2, and Hard Coal new, with each sub-category having different operational characteristics.
@@
Appendix B Example of Non-Decreasing Redispatch Costs
The following hypothetical example shows that a configuration with $k$ zones does not necessarily lead to lower redispatch costs compared to a configuration with less than $k$ zones.
Example 1.
Consider an electricity market with sellers $S=\{s_{1},s_{2},s_{3},s_{4}\}$, a buyer $b_{1}$ and an underlying transmission network $G=(V,E)$ with nodes $V=\{v_{1},v_{2},v_{3},v_{4},v_{5},v_{6}\}$ and lines $E=\{\{v_{3},v_{5}\},\{v_{4},v_{5}\}\}$. The line capacities are 50 MW for $\{v_{3},v_{5}\}$ and 100 MW for $\{v_{4},v_{5}\}$, and the two lines have the same susceptance value. The buyer $b_{1}$ is located in node $v_{5}$ and has an inelastic demand of 100 MWh. Table 14 shows, for each seller, the node in which they are located and their generation cost. Assume that each seller can generate 200 MWh. For this electricity market, consider the two zonal configurations described in Table 15.
Seller Node Cost (EUR/MWh)
$s_{1}$ $v_{1}$ 1
$s_{2}$ $v_{2}$ 3
$s_{3}$ $v_{3}$ 2
$s_{4}$ $v_{4}$ 40
Table 14: Example - Characteristics of the sellers
Configuration Zone 1 Zone 2 Zone 3 Generation Redispatch
2 Zones $\{v_{1},v_{2},v_{3},v_{4}\}$$\{v_{5},v_{6}\}$ 100 2,200
3 Zones $\{v_{2},v_{4},v_{5}\}$$\{v_{1},v_{3}\}$$\{v_{6}\}$ 200 2,300
Table 15: Example - Zonal configurations. Generation and redispatch costs are expressed in EUR.
In the 2-zone configuration, the cross-zonal capacity is equal to 150 MW. In zonal market clearing, seller $s_{1}$ generates 100 MWh to fulfill the demand of $b_{1}$. This solution is not physically feasible. The units generated by $s_{1}$ cannot be delivered to $b_{1}$ since $v_{1}$ and $v_{5}$ are not connected. A feasible solution is obtained using the minimum-cost redispatch method described in Section 2.2.2. In this solution, $s_{3}$ and $s_{4}$ each produce 50 MWh, resulting in redispatch costs equal to $1\cdot 100+2\cdot 50+40\cdot 50=2,200$ EUR.
In the 3-zone configuration, the capacity between zones 1 and 2 is 50 MW. In zonal market clearing, sellers $s_{1}$ and $s_{2}$ generate 50 MWh each. This solution is not physically feasible, since $v_{1}$ and $v_{2}$ are not connected to $v_{5}$, hence the units generated by $s_{1}$ and $s_{2}$ cannot be delivered to $b_{1}$. The redispatch method computes a feasible solution in which $s_{3}$ and $s_{4}$ produce 50 MWh each. The redispatch costs are $1\cdot 50+2\cdot 50+3\cdot 50+40\cdot 50=2,300$ EUR.

### v6

@@ 4 Conclusion
The discussion surrounding zonal and nodal pricing has a long history. The European electricity market has employed zonal pricing for many years, often with large, nationwide price zones. However, the ongoing energy transition requires a reevaluation of price zones. Based on recent data released by ENTSO-E in the context of the EU Bidding Zone Review process, we compare the short-run effects of various zonal and nodal pricing rules for the German power market. Our findings indicate that, in terms of system costs and price signals, a nodal allocation rule with non-uniform pricing leads to the lowest total costs. Costly redispatch is mitigated, and the average prices increase only slightly. Additionally, pricing rules such as Join pricing require very low side-payments, while efficiently signaling transmission bottlenecks. In contrast, the proposed zonal configurations only differ marginally regarding system costs and the differences in average prices across zones are low. Also, the effect of zonal splits on price standard deviations is low, and we do not find evidence for significant reductions in redispatch costs resulting from such splits. The study does not say that there cannot be congestion on cross-zonal lines leading to significant price differences between zones at certain times. However, in the BZR dataset that was intended to inform ACER based on realistic demand and supply scenarios for the target year 2025, we do not find significant price differences. Fears of substantial price differences after a zonal split might be overrated based on this comprehensive and up-to-date dataset. Importantly, we could provide a fair comparison with nodal pricing rules and show that the welfare gains are around 5-6%, depending on the zonal configuration.
Appendix A Market Clearing Problem - Notation
Parameters
$B,S,N,T$ Sets of buyers, sellers, network nodes, periods
$n^{*}$ Reference node
$z^{*}$ Reference zone
$n_{i}$ $B\cup S\rightarrow N$ Node in which agent $i\in B\cup S$ is located
$g_{s}$ [EUR/MWh] Marginal cost of seller $s\in S$
$h_{s}$ [EUR/MWh] No-load cost of seller $s\in S$
$\underline{R}_{s}$ Minimum uptime of seller $s\in S$ (periods)
$\underline{P}_{st},\overline{P_{st}}$ [MWh] Minimum and maximum output of seller $s\in S$ in period $t\in T$
$P_{bt}$ [MWh] Demand of buyer $b\in B$ in period $t\in T$
$\underline{F}_{nm},\overline{F}_{nm}$ [MWh] Minimum and maximum flow on the line connecting $n,m\in N$
$B_{nm}$ [pu] Susceptance of the line connecting $n,m\in N$
Variables
$x_{b,nt}$ [MW] Total consumption of buyer $b\in B$ in node $n\in N$, period $t\in T$
$y_{s,nt}\geq 0$ [MW] Total production of seller $s\in S$ in node $n\in N$, period $t\in T$
$u_{st}\in\{0,1\}$ Commitment of seller $s\in S$ in $t\in T$
$\phi_{st}\geq 0$ Startup indicator of seller $s\in S$ in period $t\in T$
$\theta_{nt}$ Voltage angle for node $n\in N$ and period $t\in T$
Table 12: Notation
Appendix B Bidding Zone Review Data
Let us discuss the data in more depth. The topology of the grid, defined by its nodes and connections, dictates the paths electricity takes from generation to consumption, with substations facilitating the necessary voltage transformations along the way. Substations act as critical nodes in the network that manage voltage levels for efficient power flow. A substation might handle multiple voltage levels. For example, a transmission substation might receive power at a very high voltage (e.g., 220kV or higher) and step it down to a lower high voltage (e.g., 110kV) for further transmission or to a medium voltage (e.g., 10-35kV) for distribution to urban or industrial areas.
The ENTSO-E dataset provides a model of Germany’s power grid, detailing 834 substations. Each of these substations encompasses one or more voltage levels, summing up to 1697 voltage levels at substations in the dataset. These voltage levels cover the spectrum from transmission ($\geq$ 220kV) to distribution ($\leq$ 110kV) and are composed of several topological nodes. Altogether, there are 2898 topological nodes listed in the data sets. Additionally, the dataset includes details for 100 substations located in the neighboring countries, which have been integrated into our analysis. For our study, we manually determined the geographical locations of all substations and computed prices for all voltage levels on these substations. After some pre-processing, which involved removing disconnected nodes, breakers, and disconnectors, our German grid model consists of 1670 nodes and 2407 lines (Schmitt, , 2023). Technically, supply and demand on different voltage levels of substations do not necessarily need to be the same. However, it turned out that prices computed for voltage levels on a substation were identical or very close in the ENTSO-E dataset, which is also the case for the nodal prices provided by ENTSO-E. Overall, the average prices that we computed were not significantly different from those in the BZR dataset for substations. There were differences on individual nodes, which is probably due to the fact that the prices computed by ENTSO-E are based on the duals of a linear programming relaxation, while we use the mixed integer linear program and pricing rules as they are defined above.
As the maximum power of each seller is given as a fixed quantity (denoting its nominal capacity), we need to manually incorporate the variability of renewable energy sources. We thus distribute the aggregate dispatched energy in each hour equally to all plants of the same type, proportional to their nominal capacities, and set the maximum power to this value. The underlying assumptions are that renewables are always dispatched at maximum power and that meteorological conditions are equal across Germany. Because hydro power plants are dispatchable, they are excluded from this logic. Out of 4537 generation units that we consider, 1709 ($\approx$37%) are subject to unit-commitment rigidities. These units have minimum uptime and minimum operating power requirements greater than 0. The data also do not include information on variable or fixed costs for renewables, which we derived from literature (Kost et al., , 2021; IRENA, , 2022; ENTSO-E, , 2022). For thermal units, variable cost information are provided, while fixed cost information were again obtained from literature (Kost et al., , 2021; Schröder et al., , 2013; Bundesnetzagentur, , 2022). We refer the reader to B.1 for more details.
An additional challenge is the mapping of generators to corresponding network nodes. The grid data specify the broad category of generating units at each node (e.g., thermal, hydro, external injections) and their nominal capacities. While renewable resources could be mapped to the seller data by matching nominal capacities, the grid data lacks additional information regarding thermal plant types. We assigned thermal generators in the grid to their operational characteristics (such as minimum uptime constraints) (Schmitt, , 2023). Firstly, we match identifiers with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur to obtain each generator’s broad thermal generation type (e.g., lignite, gas, etc.). As nominal capacities do not match between grid and seller data, and seller data have a higher granularity (e.g., Gas CCGT, Gas OCGT, etc.), we solve an integer program to assign each entity in the grid model to exactly one corresponding seller of the same plant type, with the objective to minimize the aggregated differences in capacities and number of units per plant type. We provide this integer program and the outcome of the generation matching in B.2. While this assignment might not be exact, we argue that smaller mistakes in the assignment of generators to nodes have little impact on the more general questions raised in our paper that address the benefits of different zonal and nodal configurations. The data are accessible through the ENTSO-E website.
Importantly, the data we use in our study reflects the target year 2025, while the fraction of the generator types is expected to change as Germany continues its energy transition (Energiewende) toward a greener power supply. According to the (Federal Ministry of Education and Research, , 2024), renewable energy sources are projected to account for up to 80% of gross electricity consumption by 2050. In addition, coal-fired power plants are set to be completely phased out by 2038 (The Federal Government, , 2024), while the share of combined-cycle gas turbine (CCGT) plants is expected to increase temporarily to compensate for the reduction in coal and to support the intermittent nature of renewable energy sources.
B.1 Generation Data
In what follows we outline how we derive the costs and capacities for the renewable (RES) and non-renewable (Non-RES) energy sources (Schmitt, , 2023). Since the procedures used to establish these costs differ between RES and Non-RES, we discuss them independently.
B.1.1 Renewables
To derive the costs for RES, we consider the operating costs (OPEX), fuel and CO2 costs from Table 13, which are based on (Kost et al., , 2021):
RES OPEX var Fuel CO2 OPEX fix
[EUR/kWh] [EUR/kWh] [EUR/kWh] [EUR/kW/year]
Offshore Wind 0.008 0 0 70
Onshore Wind 0.008 0 0 20
Biomass 0.004 0.04286 0.0026812 160
Photovoltaic 0 0 0 13.3
Hydro 0 0 0 36.64
Table 13: Costs for RES
We assume that RES require high initial investment costs but relatively low fixed costs. This assumption is attributed to the fact that particularly wind and solar power plants do not require ongoing expenses for fuel purchase and transportation, and have minimal maintenance costs due to the absence of moving parts. Therefore, the fixed cost $h_{s}$ is set to 0 for any renewable energy source $s\in S$. However, the fixed OPEX cannot be disregarded and are instead incorporated into the variable costs $g_{st}$.
We derive the full load hours (FLH) from Table 14 considering the generated power per production type for the year 2009 and the sum of capacities that are included in the BZR dataset.
RES FLH [h]
Offshore Wind 3,657.517
Onshore Wind 1,864.482
Biomass 5,756.5427
Photovoltaic 986.3497
Hydro 4,262.101782
Table 14: FLH for RES
Using the FLH of each production type, we determine the variable cost $c_{st}$ of a renewable unit $s\in S$ in a period $t\in T$ as follows:
$c_{st}=\text{OPEX}_{\text{var}_{f(s)}}+\text{Fuel}_{f(s)}+{\text{CO}_{2}}_{f(s)}+\frac{\text{OPEX}_{\text{fix}_{f(s)}}}{\text{FLH}_{f(s)}}$
Note that we make the simplifying assumption that variable costs are constant for all time periods. Furthermore, while allocating fixed to variable costs, we do not consider the specific fixed cost of each plant, which would lead to different variable costs for each plant. Instead, as we use the capacity and generated power per plant type, we derive fixed costs for an average plant per plant type. Table 15 shows the variable costs for the RES considered in this study. Note that in a final step, these values are anonymized for each power plant by drawing the final costs from a normal distribution, where the calculated variable costs serve as the mean and 1% of these variable costs serve as the standard deviation.
RES Variable Costs [EUR/MWh]
Offshore Wind 27.24
Onshore Wind 18.73
Biomass 77.34
Photovoltaic 13.84
Hydro 8.60
Table 15: Variable Costs for RES
The capacity of each RES unit is included in the BZR dataset as a static attribute. This static attribute does not adequately reflect the dynamic nature of RES outputs, which are influenced by weather and time of day. To calculate maximum output, we first determine the total capacity for each RES type and derive a score that represents the ratio of aggregated hourly output to total capacity. The maximum output per period for each RES unit is obtained by multiplying their capacity by this score.
B.1.2 Non-Renewables
The BZR dataset includes the variable costs for all thermal plant types. However, the fixed costs are not available, so we use the costs from (Kost et al., , 2021), which are shown in Table 16:
Technology Variable Costs [EUR/MWh]
Hard Coal 22
Lignite 32
CCGT 20
Table 16: Fixed Costs for Non-RES
Regarding the units that generate electricity from light oil, we refer to an older publication by Schröder et al., (2013), as more recent publications on the cost of electricity do not include power plants that generate electricity from oil. Concerning oil, Schröder et al., (2013) state fixed costs of 6 EUR/kW/year for the year 2010.
Based on these costs for the different technologies, we estimate yearly fixed costs for each generation unit. However, as fixed costs in our market clearing models are incurred during operating hours, these costs must be distributed over those hours. We approximate these costs using the FLH, which relate a power plant’s generated power to its capacity and represent the number of hours a unit operates at full capacity over a year. Kost et al., (2021) specify the FLH for the year 2020 shown in Table 17.
FLH [h]
Technology Lower Bound Upper Bound
Hard Coal 2,600 6,200
Lignite 5,300 7,300
Gas Turbine 500 3,000
CCGT 3,000 8,000
Table 17: FLH for Non-RES
Note that we assume that any power plant that generates electricity from gas and that is not categorized as a CCGT power plant is associated with the gas turbine technology. Furthermore, since neither Kost et al., (2021) nor Schröder et al., (2013) provide information on FLH for oil power plants, we take the estimates from a recent monitoring report of the Federal Network Agency. [^19] https://data.bundesnetzagentur.de/Bundesnetzagentur/SharedDocs/Mediathek/Monitoringberichte/monitoringbericht_energie2021.pdf In particular, we consider the installed generation capacity of 4.8 GW in 2020 and the net power generation of 4.3 TWh in 2020 implying a value of 895.83 FLH for an average power plant that produces energy from light oil. Therefore, we consider the values depicted in Table 18 for each of the plant types.
OPEX fix FLH Lower Bound FLH Upper Bound
Fuel/Plant Type [EUR/kW/year] [h] [h]
Hard Coal/old 1 22 2600 6200
Hard Coal/old 2 22 2600 6200
Hard Coal/new 22 2600 6200
Lignite/old 1 32 5300 7300
Lignite/old 2 32 5300 7300
Lignite/new 32 5300 7300
Gas/Conventional old 1 20 500 3000
Gas/Conventional old 2 20 500 3000
Gas/CCGT old 1 20 3000 8000
Gas/CCGT old 2 20 3000 8000
Gas/CCGT new 20 3000 8000
Gas/OCGT old 20 500 3000
Gas/OCGT new 20 500 3000
Gas/CCGT present 1 20 3000 8000
Gas/CCGT present 2 20 3000 8000
Light oil/- 6 895.83 895.83
Table 18: FLH for Non-RES
As the FLH represent a lower bound for the actual operating hours, we assume that the upper bound describes the actual operating hours. Thus, we denote the upper bound of the FLH of a plant technology $k$ shown in Table 18 as $\overline{FLH}_{k}$ and define the fixed cost of a non-renewable generation unit $s\in S$ as:
$h_{s}=\frac{\underline{P}_{s}\text{OPEXfix}_{k(s)}}{\overline{FLH}_{k(s)}}$
where $k(s)$ denotes the fuel/plant type of generation unit $s$. B.2 describes what non-RES capacities we assume.
B.2 Matching of Generators to Grid Locations
@@
As discussed in Section 2.2, we first identify each thermal unit’s broad plant type (e.g., hard coal, gas, etc.) by matching IDs with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur. Next, we need to distribute these plants to the sub-categories provided in the seller data. For example, all hard coal units must be distributed between Hard Coal old1, Hard Coal old2, and Hard Coal new, with each sub-category having different operational characteristics.
@@
Appendix C Analysis of Representative Days
The generation costs for each of these days are presented in Table 20. On 2009/11/23, characterized by warm and windy conditions with low demand, generation costs were minimal, and no congestion occurred in the zonal models, resulting in identical allocations for the national, 2 Zones (k) and 3 Zones configurations. In contrast, February 18, as a cold and cloudy day, saw substantially higher generation costs, but cross-zonal flow constraints had little impact on the allocation. Across all days, the shift from zonal to nodal pricing led to a similar increase in generation costs.
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 26,673.75 27,921.59 26,673.75 26,673.75 26,774.86 29,596.10
2009/03/10 31,501.65 31,512.26 31,491.45 31,667.54 31,667.18 35,200.69
2009/02/18 46,287.24 46,296.34 46,307.95 46,281.61 46,432.04 49,303.25
Table 20: Generation Costs: Nov 23, Mar 10, Feb 18
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 5,004.47 4,998.43 5,012.4 5,014.47 4,753.79 0
2009/03/10 6,295.56 6,262.94 6,305.71 6,129.62 6,129.98 0
2009/02/18 3,603.04 3,593.5 3,602.7 3,602.93 3,458.41 0
Table 21: RD-Min-Cost Redispatch Costs: Nov 23, Mar 10, Feb 18
in kEUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
2009/11/23 31,678.22 32,920.02 31,686.15 31,688.22 31,528.65 29,596.10
2009/03/10 37,797.21 37,775.2 37,797.16 37,797.16 37,797.16 35,200.69
2009/02/18 49,890.28 49,889.84 49,910.65 49,884.54 49,890.45 49,303.25
Table 22: Total Costs: Nov 25, Mar 10, Feb 18
(a) 2009/11/23 - Avg IP Prices
(b) 2009/11/23 - Avg CH Prices
(c) 2009/11/23 - Avg Join Prices
(d) 2009/03/10 - Avg IP Prices
(e) 2009/03/10 - Avg CH Prices
(f) 2009/03/10 - Avg Join Prices
(g) 2009/02/18 - Avg IP Prices
(h) 2009/02/18 - Avg CH Prices
(i) 2009/02/18 - Avg Join Prices
Figure 7: Average Nodal Prices
Table 8 summarizes the GLOCs, LLOCs, and MWPs for each selected day. On the low-demand day of 2009/11/23, a Walrasian equilibrium as in Definiton 1 is attainable for the national, 2 Zones (k) and 3 Zones models, resulting in zero GLOCs, LLOC, and MWPs. Crucially, all three pricing rules successfully obtain these equilibrium prices. The Walrasian equilibrium, however, can only persist before redispatch is considered and ceases to exist when nodal transmission constraints are taken into account.
GLOCs, LLOCs, and MWPs do not necessarily increase with average price levels. For nodal pricing, they remain relatively consistent across all days. There is a substantial increase for zonal pricing on the medium-price day of 2009/03/10, compared to the high-price day of 2009/02/18. This indicates that incentives to deviate hinge more on the quality of the price signal than the actual price level.
The simplification inherent in zonal models, which leads to redispatch, does not necessarily imply better economic properties of prices. For example, on 2009/02/18, zonal IP and Join prices result in higher GLOCs than nodal pricing. Specifically, a few zonal prices are insufficient to incentivize many market participants to adhere to the allocation.
As discussed above, CH prices always minimize GLOCs, and IP prices always yield zero LLOCs. Although GLOCs are reduced by CH pricing compared to both IP and Join pricing, their LLOCs, and often MWPs, experience a slight increase. In practical terms, this implies that congestion signals are more distorted and higher side-payments are required. In contrast, the Join pricing rule achieves a favorable trade-off between LLOCs and MWPs, requiring only slightly higher penalties (corresponding to GLOCs) than IP pricing.
in kEUR IP CH Join
National GLOC 0 0 0
LLOC 0 0 0
MWP 0 0 0
2 Zones (k) GLOC 0 0 0
LLOC 0 0 0
MWP 0 0 0
2 Zones (s) GLOC 1,774.99 1,167.39 1,774.99
LLOC 0 515.51 0
MWP 0 117.40 0
3 Zones GLOC 0 0 0
LLOC 0 0 0
MWP 0 0 0
4 Zones GLOC 1,824.29 791.11 1,824.29
LLOC 0 292.25 0
MWP 0 118.68 0
Nodal GLOC 1,290.22 273.99 1,507.04
LLOC 0 149.17 27.31
MWP 167.98 96.80 34.41
(a) GLOC, LLOC, MWP for 2009/11/23
in kEUR IP CH Join
National GLOC 2,274.12 52.96 924.26
LLOC 0 16.61 1.93
MWP 134.79 13.33 0
2 Zones (k) GLOC 3,539.04 45.62 3,539.04
LLOC 0 8.64 0
MWP 164.79 4.60 164.79
2 Zones (s) GLOC 4,730.19 260.35 918.13
LLOC 0 27.39 0.98
MWP 84.07 26.24 1.01
3 Zones GLOC 3,872.84 213.04 3,872.84
LLOC 0 172.13 0
MWP 0 153.07 0
4 Zones GLOC 787.65 751.28 970.31
LLOC 0 101.38 5.95
MWP 109.6 132.42 3.26
Nodal GLOC 1,906.47 256.65 1.942.01
LLOC 0 158.07 34.79
MWP 191.19 84.14 20.73
(b) GLOC, LLOC, MWP for 2009/03/10
in kEUR IP CH Join
National GLOC 4,442.86 318.73 2,618.64
LLOC 0 29.04 9.31
MWP 41.66 27.1 8.59
2 Zones (k) GLOC 2,258.24 502.83 2,600.59
LLOC 0 166.68 6.14
MWP 35.23 155.67 6.02
2 Zones (s) GLOC 1,727.29 319.07 1,956.49
LLOC 0 16.06 8.01
MWP 26.55 23.60 13.90
3 Zones GLOC 3,284.92 318.84 3,284.92
LLOC 0 28.68 0
MWP 46.26 26.01 46.26
4 Zones GLOC 1,523.92 464.18 2,642.54
LLOC 0 136.24 6.03
MWP 35.19 102.14 5.92
Nodal GLOC 551.27 241.74 624.76
LLOC 0 51.26 25.56
MWP 73.43 27.55 20.02
(c) GLOC, LLOC, MWP for 2009/02/18
Figure 8: Daily GLOCs, LLOCs, MWPs
Appendix D Median prices
Median prices are shown in Table 23.
in EUR/MWh National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 48.66 48.44 48.84 48.23 45.79 48.66
CH 31.09 31.17 31.27 31.11 38.53 49.41
Join 48.68 49.37 48.60 48.31 49.27 49.19
Euphemia 48.84
Table 23: Median Prices
Appendix E Modelling Differences Compared to the BZR Study
In this section, we highlight the main modelling differences between our study and the BZR study. The BZR study includes (1) the LMP study (ENTSO-E, , 2022), in which locational marginal prices (LMPs) were computed for each node in the CE grid, and (2) an assessment conducted at the CE level of the alternative bidding zone configurations identified based on the computed LMPs (ENTSO-E, , 2025).
E.1 Nodal Prices
In the LMP study, nodal prices were derived as shadow prices by solving a linear unit commitment problem, formulated as a linear program comprising Continental Europe and Ireland. In contrast, we computed nodal prices based on a mixed-integer linear problem (see A). We note that the average nodal prices between the two studies are very close: for the climate year 2009, the average value of the prices in the LMP study is 46.83 EUR/MWh, while the average value of the prices computed in this study with IP pricing is 46.63 EUR/MWh.
E.2 Zonal Market Clearing Problem
Several modelling differences relate to the zonal market clearing problem. To assess the impact of an alternative configuration at the CE level, the BZR study employs a modelling chain comprising five modules: (1) base case creation, (2) capacity calculation, (3) market coupling, (4) Operational Security Analysis (OSA) and Remedial Actions Optimisation (RAO), and (5) loop flow analysis. In module (1), a market simulation is performed for the entire EU with the goal of obtaining a market result forecast to be used for the capacity calculation (module (2)). In module (2), a flow-based market coupling and a coordinated net transfer capacity (cNTC) approach are implemented to compute the cross-zonal capacity. These approaches rely on multiple steps, such as calculating generation shift keys, zonal PTDFs and RAMs. Module (3) uses the FB parameters and cNTC values, as well as market data to perform a flow-based market coupling simulation for the CE region. The outcome of this module is the market dispatch, which is used, together with the available remedial actions (or, the available redispatch potential) as input in the OSA/RAO module. This module employs a DC load flow to identify congestions, which are subsequently alleviated in a cost-optimal way by solving a linear optimization problem. The last module receives the final dispatch and applies a power flow coloring approach to identify the share of flows not induced by cross-zonal trade (i.e., the loop flows). The implementation details of each module are not publicly available.
Modules (3) and (4) are particularly relevant for our study. Module (3) returns the hourly dispatch for each power plant, the hourly market clearing (zonal) prices, the hourly net positions of each BZ and the total socio-economic welfare from the market dispatch. While the exact price computation methodology is not detailed, zonal prices and dispatch appear to be co-optimized. This is different from the zonal market clearing problem that we consider. Our one-stage zonal market clearing problem (see Section 2.1.2) sets the cross-zonal capacity using an NTC approach. The result of this problem is the hourly dispatch for each power plant, the flows on the interconnectors and the total welfare. In a subsequent step, zonal prices are computed as described in Section 2.1.3. We also emphasize that in the BZR study, the socio-economic welfare is defined as the sum of consumer surplus, producer surplus and congestion rent. Consumer surplus is calculated as the product of the BZ load multiplied by the difference between the value of loss load [^20] The value of lost load is the estimated amount that customers receiving electricity with firm contracts would be willing to pay to avoid a disruption in their electricity service. This value is not published in (ENTSO-E, , 2025). and market clearing price. Producer surplus is calculated as the product of the generation of each plant in each BZ and the difference between the market clearing price and the marginal cost of that plant. Congestion rent is calculated as the product of price differences across each bidding zone border multiplied by the respective border flow. In contrast, in our study, the welfare is defined as the negative generation costs because load is assumed to be inelastic. The generation costs are the objective value of the allocation problem.
E.3 Redispatch Computation
Several modelling differences are linked to module (4), which focuses on redispatch computation. This module calculates the redispatch volumes with an algorithm that aims to “to remove grid congestions under full coordination in CE by minimising the necessary operational actions, aiming to reduce the resulting redispatching volumes(ENTSO-E, , 2025). Redispatch costs are computed based on the redispatch volumes considering the following components: (1) the short-term marginal costs for upward and downward redispatching using redispatching markups (see (ACER, , 2025)), (2) the costs of ensuring the availability of units for redispatching purposes, (3) the costs of grid reserves, (4) the activation costs of explicit demand side response, and (5) the start-up costs of generation units. As described in Section 2.1.2, we consider a minimum-cost redispatch approach formulated as a mixed-integer program that takes as input the zonal dispatch and modifies it such that it satisfies the DC power flow equations. [^21] This means that, unlike module (4) of the BZR study, we do not explicitly identify congested intra-zonal lines. The method assumes that all generators are redispatchable, and that redispatch is determined through a system-wide optimization, i.e., there is a cross-border cooperation between German TSOs. Moreover, we do not consider demand-side response. [^22] As described in (ENTSO-E, , 2022), demand-side response (DSR) is either explicit or implicit. Explicit DSR can participate in the redispatch process. Since our analysis is restricted to Germany, cheaper redispatch potential from other countries cannot be used to relieve a bottleneck. Our redispatch costs include the costs for upward and downward redispatch (without markups), and the start-up costs of generation units, corresponding primarily to components (1) and (5) of the BZR study. We also assume that the redispatch costs per MWh for each unit are identical to the individual costs used in the market clearing problem.
Appendix F Comparison with BZR Results
In what follows, we compare our main results against the key findings from the BZR study (ENTSO-E, , 2025). Importantly, the BZR study evaluates the alternative configurations at the CE level, whereas our study evaluates these configurations at the national German level. Also, our analysis is restricted to climate year 2009. The key modelling differences between our study and the BZR study are discussed in E.
First, the BZR report shows that all alternative configurations proposed for Germany have economic efficiency gains [^23] Economic efficiency is defined as the change in socio-economic welfare, i.e., the average change (over all climate years) in socio-economic welfare coming from the market dispatch minus the average change (over all climate years) in total additional redispatch costs at the CE level. ranging between EUR 251 - 339 million, which represent less than 1% of the simulated costs in CE. The highest gain is associated with a five-zone split. Importantly, the economic efficiency gain is not monotonically increasing with the number of zones – the configuration with three zones has an efficiency gain of EUR 251 million, while a two-zone configuration has an efficiency gain of EUR 264 million. In contrast, we show that the decrease of nodal costs relative to the total costs associated with the alternative configurations is between 7 and 8%.
Second, in the BZR results, all configurations are associated with average redispatch cost savings of $\approx$ 50% compared to the status quo BZ layout. The lowest savings ($\approx$ 30%) are observed for the climate year 2009. Our study reveals that the redispatch cost savings are below 10%, regardless of the considered alternative configuration. This discrepancy can be explained by differences in the redispatch algorithms used in the two studies (see Sections 2.1.2 and E). While the exact redispatch methodology is not fully specified in (ENTSO-E, , 2025), it assumes a fully coordinated process across Central Europe that includes demand response, battery storage, and pumped hydro storage – features not captured in our model. Additionally, redispatch costs in the BZR study account for multiple components beyond the upward, downward, and curtailment costs considered in our analysis. Similar to the results in the BZR study, our results show that the redispatch savings do not necessary increase if we consider a configuration with more zones.
Third, the BZR results show that all configurations induce two zones in Germany where average market clearing prices change in opposite directions. The northern zone(s) exhibit(s) lower prices compared to the southern one(s). For instance, in a two-zone configuration, the average price difference across all three climate years is 6.45 EUR/MWh. However, as illustrated in (ENTSO-E, , 2024), the price difference is the lowest for climate year 2009 (4.77 EUR/MWh). In contrast, our study finds that zonal price differences are below 5 EUR/MWh (see Table 6). This lower price spread can be attributed to the difference in price computation methodology and to the fact that Germany is modeled in isolation from the rest of the European bidding zones. As illustrated in (ENTSO-E, , 2025), zonal prices across bidding zones are correlated – flows and trades in one zone influence prices in others.
Finally, we calculated the generation and redispatch costs for the five-zone configuration proposed by the German TSOs following the publication of the ACER configurations. Average generation and redispatch costs were 32,220.77 kEUR and 6,014.25 kEUR, respectively, totaling 38,235.02 kEUR – this is comparable to the costs corresponding to the other zonal configurations shown in Table 2. Average zonal price differences were below 1 EUR/MWh.